# Decision Quotient: A Regime-Sensitive Complexity Theory of Exact Relevance Certification

### 16 Mar 2026 v1

@@ 2.4 Computational Model and Input Encoding
  • a finite action set $A$ given explicitly,
@@
The solver is given oracle access to $\text{Opt}(s)$ or to the utility values needed to reconstruct it. Complexity is then measured by query count, optionally paired with per-query evaluation cost. This regime separates access obstruction from full-table availability and from succinct circuit input length.
Remark 2.19 (Decision Questions vs Decision Problems).
The predicate “is coordinate set $I$ sufficient for $\mathcal{D}$?” is an encoding-independent mathematical question. A decision problem arises only after fixing how $\mathcal{D}$ and $I$ are represented. Complexity classes are properties of these encoded problems, not of the abstract predicate in isolation.
Remark 2.20 (Later Stochastic Comparison).
In the stochastic row, the preservation predicate compares the optimizer seen under coarse conditioning on $I$ with the fully conditioned optimizer. Under full support, the later bridge package reconnects this stochastic notion to the same quotient picture as the static regime.
3 Static Regime Complexity
This section studies the static regime, where the input is a finite coordinate-structured decision problem under the succinct encoding of Section 2.4. This is the baseline certification setting: no stochastic averaging, no temporal dynamics, only the exact question of whether a candidate coordinate set preserves the optimizer map. The main contribution here is structural: sufficiency is exactly relevance containment. Once that characterization is in place, the complexity classification for Minimum-Sufficient-Set follows immediately. The only genuinely higher-level static query is the anchor variant, where an existentially chosen fiber must satisfy a universal constancy condition.
3.1 Problem Definitions
Problem 3.1 (SUFFICIENCY-CHECK).
@@ 3.1 Problem Definitions
Question: Is $I$ sufficient for $\mathcal{D}$?
Problem 3.2 (MINIMUM-SUFFICIENT-SET).
@@
Question: Does there exist a sufficient set $I$ with $|I|\leq k$?
Problem 3.3 (ANCHOR-SUFFICIENCY).
Input: Decision problem $\mathcal{D}=(A,X_{1},\ldots,X_{n},U)$ and fixed coordinate set $I$.
Question: Does there exist an assignment $\alpha$ to $I$ such that $\text{Opt}(s)$ is constant on the fiber $\{s:s_{I}=\alpha\}$?
3.2 Apparent $\Sigma_{2}^{P}$ Structure
@@
$\exists I\;\forall s,s^{\prime}\in S:\quad s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime}).$
That syntax suggests a $\Sigma_{2}^{P}$ classification. The algorithm suggested by this structure is: guess a coordinate set $I$, then universally verify that every agreeing state pair induces the same optimal-action set. On its face, static exact minimization therefore looks like a genuine existential–universal search problem.
3.3 Structural Characterization of Minimum Sufficiency
Proposition 2.9 removes the outer search layer: a set is sufficient exactly when it contains every relevant coordinate, so the minimum sufficient set is simply the relevant-coordinate set. This is the central structural theorem of the static regime. Static exact minimization is therefore a relevance-counting problem, not a genuinely alternating search problem.
@@ 3.3 Structural Characterization of Minimum Sufficiency
3.4 Complexity Consequence: coNP-Completeness
@@
Membership is immediate from short counterexamples: a “no” witness is just a pair of states that agree on $I$ but induce different optimal-action sets. Hardness comes from reducing TAUTOLOGY to an empty-information instance whose optimizer map is constant exactly in the tautology case.
Theorem 3.4 (SUFFICIENCY-CHECK is coNP-complete).
@@
For membership, Proposition 2.6 gives a polynomial witness $(s,s^{\prime})$ with $s_{I}=s^{\prime}_{I}$ and $\text{Opt}(s)\neq\text{Opt}(s^{\prime})$ whenever $I$ is not sufficient, so the complement is in NP.
For hardness, reduce TAUTOLOGY to the empty-set instance. Given formula $\varphi(x_{1},\ldots,x_{n})$, build a two-action decision problem where one action tracks $\varphi$ and the other is a baseline action. Then $I=\emptyset$ is sufficient iff $\varphi$ is a tautology, because empty-set sufficiency is exactly constancy of Opt (Proposition 2.5). The reduction is polynomial in formula size.  $\blacksquare$
Theorem 3.5 (Static Structural Theorem and Complexity Consequence).
@@
Proposition 2.9 identifies the minimum sufficient set with the relevant-coordinate set. The question $|I^{*}|\leq k$?” therefore becomes the question whether at most $k$ coordinates are relevant, which is a coNP check over relevance witnesses. CoNP-hardness follows from the $k=0$ slice, which is exactly SUFFICIENCY-CHECK with $I=\emptyset$.  $\blacksquare$
3.5 Sigma-2-P Completeness
@@
Proof idea.
Membership follows the visible quantifier pattern: guess an anchor assignment and then verify constancy on its fiber. Hardness is obtained by encoding the existential variables of an $\exists\forall$ formula into the anchor choice and the universal variables into the residual coordinates that must be checked.
Theorem 3.6 (ANCHOR-SUFFICIENCY is Sigma-2-P-complete).
@@
$\exists\alpha\;\forall s\in S:\ (s_{I}=\alpha)\Rightarrow\text{Opt}(s)=\text{Opt}(s_{\alpha}).$
@@
3.6 Certificate-Size Lower Bound
@@
Theorem 3.7 (Witness-Checking Duality).
@@
Empty-set sufficiency asks whether Opt is constant on all $2^{n}$ states. A sound refutation-complete checker must be able to separate constant maps from maps that differ on some hidden antipodal partition. An adversary can place the first disagreement in any of $2^{n-1}$ independent pair slots; if fewer than $2^{n-1}$ slots are inspected, two instances remain indistinguishable to the checker but have opposite truth values. Therefore at least $2^{n-1}$ pair checks are necessary.  $\blacksquare$
@@
Under the explicit-state step-counting model, static sufficiency and static anchor sufficiency are also wrapped as abstract P predicates on inputs carrying a certified state budget. The artifact proves these upper bounds via counted searches with quadratic and linear state-space dependence respectively, and it also packages the static sufficiency search as an explicit correctness-plus-step-bound witness and as part of the unified finite-search summary theorem. (L: DC70-71, DC76, DC80-82.)
@@
4 Stochastic Regime Complexity
@@ 4.1 Preservation and Decisiveness
4.2 Stochastic Decision Problems
Definition 4.1 (Stochastic Decision Problem).
@@ 4.2 Stochastic Decision Problems
Definition 4.2 (Distribution-Conditional Sufficiency).
@@
$\text{Opt}^{\mathrm{stoch}}_{I}(\alpha):=\arg\max_{a\in A}\,\mathbb{E}[U(a,S)\mid S_{I}=\alpha].$
A coordinate set $I$ is stochastically sufficient if conditioning on $I$ preserves the fully conditioned optimizer:
$\forall s\in S:\quad\text{Opt}^{\mathrm{stoch}}_{I}(s_{I})=\text{Opt}(s).$
@@
The conditional optimizer $\text{Opt}^{\mathrm{stoch}}_{I}(\alpha)$ aggregates expected utilities over the entire fiber $\{s:s_{I}=\alpha\}$ using the distribution $P$. By contrast, the full-information optimizer $\text{Opt}(s)$ conditions on the exact state $s$ itself. Preservation asks whether this averaging over a fiber ever changes which action is optimal; decisiveness asks whether, after averaging, the optimal action on each fiber is uniquely determined.
Definition 4.3 (Stochastic Decisiveness).
Using the same conditional optimizer map $\text{Opt}^{\mathrm{stoch}}_{I}$, a coordinate set $I$ is stochastically decisive if every admissible fiber has a uniquely determined conditional optimal action:
$\forall\alpha\in\mathrm{supp}(S_{I})\;\exists a\in A:\quad\text{Opt}^{\mathrm{stoch}}_{I}(\alpha)=\{a\}.$
@@
Consider a stochastic decision problem with $S=\{s_{1},s_{2}\}$, $A=\{a,b\}$, uniform distribution $P(s_{1})=P(s_{2})=1/2$, and utilities $U(a,s_{1})=2$, $U(b,s_{1})=1$, $U(a,s_{2})=0$, $U(b,s_{2})=3$. For $I=\emptyset$: (i) Preservation: $\mathbb{E}[U(a,S)]=1$, $\mathbb{E}[U(b,S)]=2$, so $\text{Opt}^{\mathrm{stoch}}_{\emptyset}=\{b\}$; but $\text{Opt}(s_{1})=\{a\}$, so preservation fails. (ii) Decisiveness: The empty fiber itself has a unique conditional optimum $\{b\}$, so $I=\emptyset$ is decisive. This shows decisiveness is strictly stronger than preservation.
4.3 Stochastic Status at a Glance
Table 2.
Regime classification summary for the stochastic row. Open entries mark explicit scope boundaries. Under full support, stochastic preservation is completely classified in the full-support base case.
@@ 4.3 Stochastic Status at a Glance
Stochastic sufficiency (preservation) Polynomial-time under explicit-state encoding; bridges to static sufficiency and the optimizer quotient Explicit-state decider; full-support bridge package No succinct-encoding hardness claimed
@@
Stochastic decisiveness Polynomial-time under explicit-state encoding; PP-hard under succinct encoding Reduction core; finite witness/checking schemata No oracle-machine formalization
Minimum / anchor decisiveness PP-hard; in $\textsf{NP}^{\textsf{PP}}$ at paper level Existential witness/checking schemata; bounded explicit-state searches No oracle-class membership formalization
@@
Problem 4.4 (STOCHASTIC-SUFFICIENCY-CHECK).
@@
Question: Is $I$ stochastically sufficient?
4.4 Stochastic Sufficiency
Theorem 4.5 (Stochastic Sufficiency is Decidable in Explicit State).
@@ 4.4 Stochastic Sufficiency
For a fixed information set $I$, the verifier scans the finite state space and checks whether the conditional fiber optimizer $\text{Opt}^{\mathrm{stoch}}_{I}(s_{I})$ agrees with the fully conditioned optimizer $\text{Opt}(s)$ at every state $s$. The artifact contains a counted explicit-state search witnessing exactly this predicate together with a linear-in-$|S|$ step bound (L: DC91-93.) Hence stochastic sufficiency is polynomial-time decidable in the explicit-state model.  $\blacksquare$
@@
Under the succinct encoding of Section 2.4, Stochastic-Preservation-Check for general distributions (without the full-support assumption) is coNP-hard.
The structured justification is anchored to existing theorems and mechanized lemmas. First, what is known: under full support, preservation inherits the static coNP classification via Proposition 4.7 and Theorem 4.11; the explicit-state base query is polynomial for all distributions by Theorem 4.5; and decisiveness under the same encoding is PP-hard by Theorem 4.19.
Second, the structural obstacle: the support-sensitive obstruction lemmas in the artifact (Proposition 4.14) isolate the exact mechanism. Preservation quantifies over all states while stochastic conditioning weights by probability. Zero-probability states can therefore witness static violations that the distribution never observes; full-support transfer fails exactly when such states are permitted. This is a structural obstruction, and the mechanized artifact isolates the finite combinatorial core any reduction must respect (L: OU12, DC96, DC98).
Third, why coNP rather than PP: preservation is syntactically universal, because it asks whether the coarse optimizer matches the full optimizer at every state, so its quantifier shape matches the static coNP classification. Decisiveness is PP-hard because it requires majority-style conditional expected-utility comparisons. Preservation compares optimizer sets rather than aggregating utilities, so the hardness source is counterexample exclusion (coNP) rather than majority counting (PP); support-sensitivity complicates reduction engineering but not the high-level quantifier structure.
Reduction intuition.
A candidate hardness reduction would encode static counterexample pairs into zero-probability witness states of the distribution. Concretely, from a TAUTOLOGY instance $\varphi$, one could construct a succinct stochastic decision problem where the distribution assigns zero mass to a carefully chosen set of states that encode the $\varphi$-violating assignments. Preservation must check the universal condition over ALL states, including these zero-probability witnesses, so a coNP-style verification can detect the hidden TAUTOLOGY counterexample even though the distribution never observes it. This intuition aligns with the support-sensitive obstruction lemmas (Proposition 4.14) but requires overcoming the technical challenge of succinctly encoding witness-state support.
Open Problem 4.2. Determine the exact succinct-encoding complexity of Stochastic-Minimum-Preservation and Stochastic-Anchor-Preservation for general distributions. Under Conjecture 4.1, these inherit coNP-completeness and $\Sigma_{2}^{P}$-completeness respectively, matching the static regime; resolving the conjecture would complete the final open cell in the regime matrix (Table 2).
Proposition 4.6 (Stochastic Sufficiency Implies Static Sufficiency).
If $I$ is stochastically sufficient, then $I$ is sufficient for the underlying static decision problem.
(L: DC94.)
@@
If the coarse conditional optimizer agrees with the fully conditioned optimizer at every state, then any two states that agree on $I$ also have equal full-information optimal-action sets. So the original static optimizer already factors through the projection to $I$.  $\blacksquare$
Proposition 4.7 (Full-Support Equivalence).
Assume the distribution $P$ has full support and the action set is nonempty. Then $I$ is statically sufficient if and only if it is stochastically sufficient.
(L: DC95-96.)
@@
The forward direction uses full support to show that when the static optimizer is constant on each $I$-fiber, conditional averaging cannot change the optimizer on that fiber. The reverse direction is Proposition 4.6. Both directions are mechanized.  $\blacksquare$
Proposition 4.8 (Quotient Equivalence under Full Support).
Under the hypotheses of Proposition 4.7, the stochastic fiber quotient induced by $I$ coincides with the original decision quotient.
(L: DC97-99.)
@@
Under stochastic sufficiency, the conditional fiber optimizer and the fully conditioned optimizer agree statewise. Hence the induced stochastic equivalence relation matches the original decision-equivalence relation, so the quotient setoids coincide. Under full support, Proposition 4.7 supplies the needed preservation hypothesis.  $\blacksquare$
Problem 4.9 (STOCHASTIC-MINIMUM-PRESERVATION).
@@
Question: Does there exist a stochastically sufficient coordinate set $I$ with $|I|\leq k$?
Problem 4.10 (STOCHASTIC-ANCHOR-PRESERVATION).
@@
Question: Does there exist an anchor state $s_{0}$ such that for every state $s$ agreeing with $s_{0}$ on $I$, one has $\text{Opt}^{\mathrm{stoch}}_{I}(s_{I})=\text{Opt}(s)$?
Theorem 4.11 (Full-Support Inheritance for Preservation Variants).
@@
    For each fixed $I$, Stochastic-Anchor-Preservation is equivalent to the underlying static Anchor-Sufficiency query on $I$.
@@
For minimum preservation, Proposition 4.7 applies pointwise to each candidate coordinate set $I$, so the existential minimization problem is unchanged under full support. For anchor preservation, if one anchor fiber preserves the full-information optimizer, then the optimizer is already constant on that fiber, giving static anchor sufficiency. Conversely, if the static optimizer is constant on one $I$-fiber, then under full support the same conditional-expectation argument used in Proposition 4.7 shows that the coarse conditional optimizer agrees with the full-information optimizer throughout that fiber. The complexity claims therefore inherit directly from Theorems 3.5 and 3.6.  $\blacksquare$
Proposition 4.12 (Explicit Finite Search for Preservation Variants).
@@
Theorem 4.13 (Explicit-State Tractability for Preservation Variants).
@@
Proposition 4.14 (General-Distribution Obstructions and Support-Sensitive Bridge).
Let $I$ be a coordinate set.
@@
    If $I$ is stochastically sufficient, then $I$ contains every relevant coordinate of the underlying static decision problem.
@@
    If the underlying static decision problem is sufficient on $I$ and every $I$-fiber contains at least one positive-probability state, then $I$ is stochastically sufficient.
@@
    If the underlying static decision problem is anchor-sufficient on $I$ and the anchor fiber contains a positive-probability state, then $I$ is stochastically anchor-preserving.
@@
For stochastic sufficiency in the preservation sense, this paper proves explicit-state tractability for the base, minimum, and anchor queries (Theorem 4.5, Proposition 4.12, and Theorem 4.13), static-to-stochastic bridge theorems (Propositions 4.6 and 4.7), quotient equivalence under full support (Proposition 4.8), and full-support inheritance for the minimum and anchor variants (Theorem 4.11). Beyond full support, the paper now also isolates unconditional obstruction lemmas and positive-support bridge results (Proposition 4.14). This yields a complete classification for stochastic preservation under full support, the full-support base case in which every state has nonzero probability, together with support-sensitive partial theory beyond that base case. The general-distribution case still introduces technical complications from zero-probability fibers, and the paper does not claim a full general-distribution classification for the minimum and anchor preservation variants or a succinct-encoding hardness classification for the preservation predicate itself. The PP-hardness results in the stochastic row concern decisiveness and its anchor/minimum variants, not preservation itself.
Problem 4.15 (STOCHASTIC-DECISIVENESS-CHECK).
@@
Question: Is $I$ stochastically decisive?

### 31 Mar 2026 v2

@@ II-D Computational Model and Input Encoding
  • a finite action set AA given explicitly,
@@
The solver is given oracle access to $\mathrm{Opt}(s)$ or to the utility values needed to reconstruct it. Complexity is then measured by query count, optionally paired with per-query evaluation cost. This regime separates access obstruction from full-table availability and from succinct circuit input length.
Remark II.19 (Decision Questions vs Decision Problems).
The predicate “is coordinate set II sufficient for $\mathcal{D}$?” is an encoding-independent mathematical question. A decision problem arises only after fixing how $\mathcal{D}$ and II are represented. Complexity classes are properties of these encoded problems, not of the abstract predicate in isolation.
Remark II.20 (Later Stochastic Comparison).
In the stochastic row, the preservation predicate compares the optimizer seen under coarse conditioning on II with the fully conditioned optimizer. Under full support, the later bridge package reconnects this stochastic notion to the same quotient picture as the static regime.
III Static Regime Complexity
This section studies the static regime, where the input is a finite coordinate-structured decision problem under the succinct encoding of Section II-D. This is the baseline certification setting: no stochastic averaging, no temporal dynamics, only the exact question of whether a candidate coordinate set preserves the optimizer map. The main contribution here is structural: sufficiency is exactly relevance containment. Once that characterization is in place, the complexity classification for Minimum-Sufficient-Set follows immediately. The only genuinely higher-level static query is the anchor variant, where an existentially chosen fiber must satisfy a universal constancy condition.
III-A Problem Definitions
Problem III.1 (SUFFICIENCY-CHECK).
@@ III-A Problem Definitions
Question: Is II sufficient for $\mathcal{D}$?
Problem III.2 (MINIMUM-SUFFICIENT-SET).
@@
Question: Does there exist a sufficient set II with $|I|\leq k$?
Problem III.3 (ANCHOR-SUFFICIENCY).
Input: Decision problem $\mathcal{D}=(A,X_{1},\ldots,X_{n},U)$ and fixed coordinate set II.
Question: Does there exist an assignment $\alpha$ to II such that $\mathrm{Opt}(s)$ is constant on the fiber $\{s:s_{I}=\alpha\}$?
III-B Apparent $\Sigma_{2}^{P}$ Structure
@@
$\exists I\;\forall s,s^{\prime}\in S:\quad s_{I}=s^{\prime}_{I}\implies\mathrm{Opt}(s)=\mathrm{Opt}(s^{\prime}).$
That syntax suggests a $\Sigma_{2}^{P}$ classification. The algorithm suggested by this structure is: guess a coordinate set II, then universally verify that every agreeing state pair induces the same optimal-action set. On its face, static exact minimization therefore looks like a genuine existential–universal search problem.
III-C Structural Characterization of Minimum Sufficiency
Proposition II.9 removes the outer search layer: a set is sufficient exactly when it contains every relevant coordinate, so the minimum sufficient set is simply the relevant-coordinate set. This is the central structural theorem of the static regime. Static exact minimization is therefore a relevance-counting problem, not a genuinely alternating search problem.
@@ III-C Structural Characterization of Minimum Sufficiency
III-D Complexity Consequence: coNP-Completeness
@@
Membership is immediate from short counterexamples: a “no” witness is just a pair of states that agree on II but induce different optimal-action sets. Hardness comes from reducing TAUTOLOGY to an empty-information instance whose optimizer map is constant exactly in the tautology case.
Theorem III.4footnotemark: footnotetext: : DQ40-41 (SUFFICIENCY-CHECK is coNP-complete).
@@
For membership, Proposition II.6 gives a polynomial witness $(s,s^{\prime})$ with $s_{I}=s^{\prime}_{I}$ and $\mathrm{Opt}(s)\neq\mathrm{Opt}(s^{\prime})$ whenever II is not sufficient, so the complement is in NP.
For hardness, reduce TAUTOLOGY to the empty-set instance. Given formula $\varphi(x_{1},\ldots,x_{n})$, build a two-action decision problem where one action tracks $\varphi$ and the other is a baseline action. Then $I=\emptyset$ is sufficient iff $\varphi$ is a tautology, because empty-set sufficiency is exactly constancy of $\mathrm{Opt}$ (Proposition II.5). The reduction is polynomial in formula size.  $\blacksquare$
Theorem III.5footnotemark: footnotetext: : DQ35-36 (Static Collapse Theorem).
@@
Proposition II.9 identifies the minimum sufficient set with the relevant-coordinate set. The question $|I^{*}|\leq k$?” therefore becomes the question whether at most $k$ coordinates are relevant, which is a coNP check over relevance witnesses. CoNP-hardness follows from the $k=0$ slice, which is exactly SUFFICIENCY-CHECK with $I=\emptyset$.  $\blacksquare$
III-E Sigma-2-P Completeness
@@
Theorem III.6footnotemark: footnotetext: : DQ98, CC4 (ANCHOR-SUFFICIENCY is Sigma-2-P-complete).
@@
$\exists\alpha\;\forall s\in S:\ (s_{I}=\alpha)\Rightarrow\mathrm{Opt}(s)=\mathrm{Opt}(s_{\alpha}).$
@@
III-F Certificate-Size Lower Bound
@@
Theorem III.7footnotemark: footnotetext: : WD1-3 (Witness-Checking Duality).
@@
Empty-set sufficiency asks whether $\mathrm{Opt}$ is constant on all $2^{n}$ states. A sound refutation-complete checker must be able to separate constant maps from maps that differ on some hidden antipodal partition. An adversary can place the first disagreement in any of $2^{n-1}$ independent pair slots; if fewer than $2^{n-1}$ slots are inspected, two instances remain indistinguishable to the checker but have opposite truth values. Therefore at least $2^{n-1}$ pair checks are necessary.  $\blacksquare$
@@
Under the explicit-state step-counting model, static sufficiency and static anchor sufficiency are also wrapped as abstract P predicates on inputs carrying a certified state budget. The artifact proves these upper bounds via counted searches with quadratic and linear state-space dependence respectively, and it also packages the static sufficiency search as an explicit correctness-plus-step-bound witness and as part of the unified finite-search summary theorem. 141414: DC70-71, DC76, DC80-82.
@@
IV Stochastic Regime Complexity
@@ IV-A Preservation and Decisiveness
Example IV.1 (Preservation without Decisiveness).
Consider a decision problem with two coordinates $X_{1},X_{2}\in\{0,1\}$, actions $A=\{0,1\}$, uniform distribution over states, and utility $U(a,s)=1$ if $a=s_{1}\oplus s_{2}$ (XOR), 0 otherwise. Hiding $X_{2}$ preserves the optimizer: given any value of $X_{1}$, the conditional optimal action is still $s_{1}\oplus s_{2}$ (since the distribution is uniform, the conditional expectation equals pointwise utility). However, $X_{1}$ alone is not decisive: for each fixed $X_{1}$, both $s_{2}=0$ and $s_{2}=1$ yield different optimal actions (0 and 1, or 1 and 0), so no single action is conditionally optimal across the entire fiber.
Example IV.2 (Decisiveness without Preservation).
Consider the same coordinate structure but with utility $U(a,s)=1$ if $a=s_{1}$, 0 otherwise. Here $X_{1}$ is decisive: for each $x_{1}$, action $x_{1}$ is uniquely optimal. However, $X_{1}$ does not preserve the full optimizer: when $s_{2}$ is revealed, the optimal action depends on $s_{1}$ alone, but the utility values differ across $s_{2}$, so the conditional expectation at $X_{1}=x_{1}$ averages over utilities that $X_{1}$ cannot distinguish.
IV-B Stochastic Decision Problems
Definition IV.3 (Stochastic Decision Problem).
@@ IV-B Stochastic Decision Problems
Definition IV.4 (Distribution-Conditional Sufficiency).
@@
$\mathrm{Opt}^{\mathrm{stoch}}_{I}(\alpha):=\arg\max_{a\in A}\,\mathbb{E}[U(a,S)\mid S_{I}=\alpha].$
A coordinate set II is stochastically sufficient if conditioning on II preserves the fully conditioned optimizer:
$\forall s\in S:\quad\mathrm{Opt}^{\mathrm{stoch}}_{I}(s_{I})=\mathrm{Opt}(s).$
@@
The conditional optimizer $\mathrm{Opt}^{\mathrm{stoch}}_{I}(\alpha)$ aggregates expected utilities over the entire fiber $\{s:s_{I}=\alpha\}$ using the distribution $P$. By contrast, the full-information optimizer $\mathrm{Opt}(s)$ conditions on the exact state $s$ itself. Preservation asks whether this averaging over a fiber ever changes which action is optimal; decisiveness asks whether, after averaging, the optimal action on each fiber is uniquely determined.
Definition IV.5 (Stochastic Decisiveness).
Using the same conditional optimizer map $\mathrm{Opt}^{\mathrm{stoch}}_{I}$, a coordinate set II is stochastically decisive if every admissible fiber has a uniquely determined conditional optimal action:
$\forall\alpha\in\mathrm{supp}(S_{I})\;\exists a\in A:\quad\mathrm{Opt}^{\mathrm{stoch}}_{I}(\alpha)=\{a\}.$
@@
Consider a stochastic decision problem with $S=\{s_{1},s_{2}\}$, $A=\{a,b\}$, uniform distribution $P(s_{1})=P(s_{2})=1/2$, and utilities $U(a,s_{1})=2$, $U(b,s_{1})=1$, $U(a,s_{2})=0$, $U(b,s_{2})=3$. For $I=\emptyset$: (i) Preservation: $\mathbb{E}[U(a,S)]=1$, $\mathbb{E}[U(b,S)]=2$, so $\mathrm{Opt}^{\mathrm{stoch}}_{\emptyset}=\{b\}$; but $\mathrm{Opt}(s_{1})=\{a\}$, so preservation fails. (ii) Decisiveness: The empty fiber itself has a unique conditional optimum $\{b\}$, so $I=\emptyset$ is decisive. This shows decisiveness is strictly stronger than preservation.
IV-C Stochastic Status at a Glance
@@
Stochastic sufficiency (preservation) Polynomial-time under explicit-state encoding; bridges to static sufficiency Explicit-state decider; full-support bridge package No succinct-encoding hardness claimed
@@
Stochastic decisiveness Polynomial-time under explicit-state encoding; $P$-hard under succinct encoding Reduction core; finite witness/checking schemata No oracle-machine formalization
Minimum / anchor decisiveness $P$-hard; in $\textsf{NP}^{\textsf{PP}}$ at paper level Existential witness/checking schemata; bounded explicit-state searches No oracle-class membership formalization
TABLE II: Regime classification summary for the stochastic row. Open entries mark explicit scope boundaries. Under full support, stochastic preservation is completely classified in the full-support base case.
@@
Problem IV.6 (STOCHASTIC-SUFFICIENCY-CHECK).
@@
Question: Is II stochastically sufficient?
IV-D Stochastic Sufficiency
Theorem IV.7footnotemark: footnotetext: : DC91-93, OU12 (Stochastic Sufficiency is Decidable in Explicit State).
@@ IV-D Stochastic Sufficiency
For a fixed information set II, the verifier scans the finite state space and checks whether the conditional fiber optimizer $\mathrm{Opt}^{\mathrm{stoch}}_{I}(s_{I})$ agrees with the fully conditioned optimizer $\mathrm{Opt}(s)$ at every state $s$. The artifact contains a counted explicit-state search witnessing exactly this predicate together with a linear-in-$|S|$ step bound 161616: DC91-93. Hence stochastic sufficiency is polynomial-time decidable in the explicit-state model.  $\blacksquare$
@@
Under the succinct encoding of Section II-D, Stochastic-Preservation-Check for general distributions (without the full-support assumption) is coNP-hard.
Intuition: preservation remains syntactically universal (statewise optimizer agreement), so its quantifier profile aligns with coNP-style counterexample exclusion rather than $P$-style majority comparison. The support-sensitive obstruction lemmas (Proposition IV.16) isolate why full-support transfer fails: zero-probability states can witness violations even when they never appear in conditional averages. The mechanized finite core for this obstruction is available 171717: OU12, DC96, DC98.
Open Problem 4.2. Determine the exact succinct-encoding complexity of Stochastic-Minimum-Preservation and Stochastic-Anchor-Preservation for general distributions. Under Conjecture IV-D, these inherit coNP-completeness and $\Sigma_{2}^{P}$-completeness respectively, matching the static regime; resolving the conjecture would complete the final open cell in the regime matrix (Table 2).
Proposition IV.8footnotemark: footnotetext: : DC94 (Stochastic Sufficiency Implies Static Sufficiency).
If II is stochastically sufficient, then II is sufficient for the underlying static decision problem.
@@
If the coarse conditional optimizer agrees with the fully conditioned optimizer at every state, then any two states that agree on II also have equal full-information optimal-action sets. So the original static optimizer already factors through the projection to II.  $\blacksquare$
Proposition IV.9footnotemark: footnotetext: : DC95-96 (Full-Support Equivalence).
Assume the distribution $P$ has full support and the action set is nonempty. Then II is statically sufficient if and only if it is stochastically sufficient.
@@
The forward direction uses full support to show that when the static optimizer is constant on each II-fiber, conditional averaging cannot change the optimizer on that fiber. The reverse direction is Proposition IV.8. Both directions are mechanized.  $\blacksquare$
Proposition IV.10footnotemark: footnotetext: : DC97-99 (Quotient Equivalence under Full Support).
Under the hypotheses of Proposition IV.9, the stochastic fiber quotient induced by II coincides with the original decision quotient.
@@
Under stochastic sufficiency, the conditional fiber optimizer and the fully conditioned optimizer agree statewise. Hence the induced stochastic equivalence relation matches the original decision-equivalence relation, so the quotient setoids coincide. Under full support, Proposition IV.9 supplies the needed preservation hypothesis.  $\blacksquare$
Problem IV.11 (STOCHASTIC-MINIMUM-PRESERVATION).
@@
Question: Does there exist a stochastically sufficient coordinate set II with $|I|\leq k$?
Problem IV.12 (STOCHASTIC-ANCHOR-PRESERVATION).
@@
Question: Does there exist an anchor state $s_{0}$ such that for every state $s$ agreeing with $s_{0}$ on II, one has $\mathrm{Opt}^{\mathrm{stoch}}_{I}(s_{I})=\mathrm{Opt}(s)$?
Theorem IV.13footnotemark: footnotetext: : DQ94-95 (Full-Support Inheritance for Preservation Variants).
@@
    For each fixed II, Stochastic-Anchor-Preservation is equivalent to the underlying static Anchor-Sufficiency query on II.
@@
For minimum preservation, Proposition IV.9 applies pointwise to each candidate coordinate set II, so the existential minimization problem is unchanged under full support. For anchor preservation, if one anchor fiber preserves the full-information optimizer, then the optimizer is already constant on that fiber, giving static anchor sufficiency. Conversely, if the static optimizer is constant on one II-fiber, then under full support the same conditional-expectation argument used in Proposition IV.9 shows that the coarse conditional optimizer agrees with the full-information optimizer throughout that fiber. The complexity claims therefore inherit directly from Theorems III.5 and III.6.  $\blacksquare$
Proposition IV.14footnotemark: footnotetext: : DQ86-89 (Explicit Finite Search for Preservation Variants).
@@
Theorem IV.15footnotemark: footnotetext: : DQ92-93 (Explicit-State Tractability for Preservation Variants).
@@
Proposition IV.16footnotemark: footnotetext: : DQ90-91, DQ96-97 (General-Distribution Obstructions and Support-Sensitive Bridge).
Let II be a coordinate set.
@@
    If II is stochastically sufficient, then II contains every relevant coordinate of the underlying static decision problem.
@@
    If the underlying static decision problem is sufficient on II and every II-fiber contains at least one positive-probability state, then II is stochastically sufficient.
@@
    If the underlying static decision problem is anchor-sufficient on II and the anchor fiber contains a positive-probability state, then II is stochastically anchor-preserving.
@@
This yields a complete preservation classification under full support, together with support-sensitive partial theory beyond that base case. The $P$-hardness results in the stochastic row concern decisiveness, not preservation.
Problem IV.17 (STOCHASTIC-DECISIVENESS-CHECK).
@@
Question: Is II stochastically decisive?